# thanks: These two authors contribute equallythanks: These two authors contribute equally

### v1

@@ Abstract
We develop a graph-based method to study the entanglement entropy of Calderbank-Shor-Steane quantum codes. This method offers a straightforward interpretation for the entanglement entropy of quantum error correcting codes through graph-theoretical concepts, shedding light on the origins of both the local and long-range entanglement. Furthermore, it inspires an efficient computational scheme for evaluating the entanglement entropy. We illustrate the method by calculating the von Neumann entropy of subsystems in toric codes and a special type of quantum low-density-parity check codes, known as bivariate bicycle codes, and by comparing the scaling behavior of entanglement entropy with respect to subsystem size. Our method provides a new perspective for understanding the entanglement structure in quantum many-body systems.
@@
Introduction. Entanglement plays a crucial role in understanding the physics of quantum many-body systems, including topological matters Kitaev and Preskill (2006); Levin and Wen (2006); Eisert et al. (2010), the black hole information paradox Hawking (1976); Page (1993), information scrambling in a thermalization process Hayden and Preskill (2007), the connection between geometry and entanglement in quantum gravity Almheiri et al. (2015); Pastawski et al. (2015), and symmetry breaking/restoration Ares et al. (2023, 2024); Liu et al. (2024). In particular, the scaling law of the entanglement entropy with respect to subsystem size reveals key properties of quantum systems, e.g., discriminating critical and non-critical phases Vidal et al. (2003); Latorre et al. (2003).
Entanglement is introduced in quantum error correcting codes to encode logical information in a non-local manner and protect it from local errors Shor (1995). This is manifested in their notable feature of local indistinguishability Knill and Laflamme (1997); Bennett et al. (1996). There exists a trade-off between the amount of entanglement and the code distance Bravyi et al. (2024a); Li et al. (2024): the lower bound of the geometric entanglement is proportional to the code distance, which implies that more entanglement is required to correct more errors.
@@
Given the unique role of entanglement in quantum error correcting codes, it is important to develop methods to calculate the amount of entanglement within quantum codes. A group-theoretical method to evaluate the entanglement entropy for stabilizer codes was proposed by Refs. Fattal et al. (2004); Hamma et al. (2005a, b). It was used to study the topological entanglement entropy in toric codes Flammia et al. (2009) and other topological codes Kargarian (2008). Recently, a graph-based method was developed to study higher-rank topological phases Ebisu (2023).
@@
In this paper, we develop a different graph-based method to calculate the explicit form of the reduced density matrix and the entanglement entropy of Calderbank-Shor-Steane (CSS) quantum codes. It provides a clear picture on the origin of the local entanglement, which indicates the area law, and the long-range topological entanglement in the toric codes. Furthermore, it inspires an efficient algorithm to evaluate the entanglement entropy for all CSS codes, e.g., quantum low-density-parity-check (qLDPC) codes Tillich and Zémor (2013); Kovalev and Pryadko (2013); Breuckmann and Terhal (2016); Panteleev and Kalachev (2021); Breuckmann and Eberhardt (2021a); Panteleev and Kalachev (2022); Leverrier and Zémor (2022); Higgott and Breuckmann (2023); Wang et al. (2023); Lin and Pryadko (2024); Breuckmann and Eberhardt (2021b). We demonstrate the graph-based method by calculating the entanglement entropy of the bivariate-bicycle codes Bravyi et al. (2024b) and studying its scaling law with respect to the size of the subsystem.
Two-dimensional toric code. We briefly introduce some basic concepts in graph theory Diestel (2024) that are essential for the subsequent discussions. A graph $G(V,E)$ consists of a set of vertices $V$ and a set of edges $E$. An edge connecting two vertices $v_{i},v_{j}\in V$ is represented by the pair $(v_{i},v_{j})$.
@@
A path is a sequence of distinct vertices connected by edges, and a cycle is a path that starts and ends at the same vertex. A connected graph is one in which there exists at least one path between any pair of vertices. A tree graph is a connected graph that contains no cycles. A spanning tree of a graph $G$ is a tree subgraph that includes all the vertices of $G$.
@@
Two-dimensional toric codes can be represented by simple two-dimensional graphs, where qubits are placed on edges. The vertices and faces of the graph represent stabilizers of type $X$ and $Z$, respectively Kitaev (2003). There are two distinct sets of cycles in the graph, the contractible cycles and non-contractible cycles. The former corresponds to $Z$-type stabilizers, while the latter represents logical Pauli $Z$ operators. Since all stabilizers have an eigenvalue of one in the code subspace, the sum of qubit values along any contractible cycle is zero (mod 2), imposing a constraint on these qubit values. The sum of qubit values along a non-contractible cycle can be either 0 or 1, depending on the specific logical state. If one of the logical Pauli $Z$ operators takes a specific eigenvalue, the qubit values along the corresponding non-contractible cycle are also subject to a constraint.
@@
We are concerned with the reduced density matrix and the von Neumann entropy of a subsystem $A$, a subset of qubits, of a toric code. The remaining qubits of the toric code are in the complementary subsystem $B$, as shown in Fig. 1. There are cycles that are entirely enclosed by subsystem $A$ or subsystem $B$, as well as cycles shared by both subsystems, which are referred to as boundary stabilizers. The boundary of subsystem $A$, denoted as $\partial A$, is defined as the set of edges (qubits) that belong to the boundary stabilizers but are not part of subsystem $A$.
@@
Figure 1: Illustration of subsystems and their spanning trees in a toric code that is embedded in a square lattice with periodic boundary conditions. (a) Subsystem $A$ consists of qubits attached on blue edges and contains one cycle as indicated by the orange face. Subsystem $B$ consists of qubits that are not included in subsystem $A$. The stabilizer operators shared by both subsystems $A$ and $B$ are indicated by pink faces. The boundary of subsystem $A$ is represented by green edges. The gray arrows indicate the closed shortest walk along the boundary. (b) The spanning tree of subsystem $A$ is obtained by removing one of the four qubits (blue cross) along the orange cycle. The spanning tree of subsystem $B$ is obtained by removing appropriate qubits (black crosses) in subsystem $B$ and one qubit along the boundary (green cross). The blue vertices labelled by $\{v_{i}\}$ are the common vertices between the spanning trees of $A$ and $B$.
Consider a connected subsystem $A$ that contains only contractible cycles and its boundary divides a plane into two regions, which are connected areas surrounded by the edges of $\partial A$. The boundary of a region includes the vertices and edges that are incident to the region, and contains a single cycle, as shown in Fig. 1. The qubit values along a cycle are not independent because they are subject to a constraint. In particular, the value of one qubit is fully determined by the values of the remaining independent quibts along the cycle. This implies that taking into account the values of the independent qubits and the constraint is sufficient to determine the state of all qubits along the cycle. From the perspective of graph theory, this is equivalent to deleting one edge from the cycle and transforming it into a path. For a system that contains many cycles, one can selectively delete appropriate edges to remove all cycles, resulting in a spanning tree. The state of the system is then fully determined by the independent qubits in the spanning tree and the associated constraints.
To facilitate further discussion in the language of graph theory, we map subsystems $A$ and $B$, as well as the boundary $\partial A$, to graphs by incorporating their vertices. The symbols $A$, $B$ and $\partial A$ are used to refer to both the sets of qubits and their corresponding graphs. By deleting the appropriate edges to remove all the cycles within the graphs $A$ and $B$, we obtain spanning trees $T_{A}(V_{T_{A}},E_{T_{A}})$ and $T_{B}(V_{T_{B}},E_{T_{B}})$, respectively. Since the boundary $\partial A$ is a subgraph of $B$ and contains a single cycle, we delete one edge from the cycle. The union of the spanning trees $T_{A}$ and $T_{B}$ forms joint cycles, introducing new constraints. This implies that the state of subsystem $A$ depends on the state of subsystem $B$, and vice versa.
The overlap between the set of vertices of the spanning tree $T_{A}$ and that of the boundary $\partial A$ is denoted as $V_{T_{A}}\bigcap V_{\partial A}=\left\{v_{1},....,v_{m}\right\}$, where $m$ represents the total number of common vertices. These vertices are ordered counterclockwise along the shortest closed walk on the boundary, with $v_{1}$ chosen arbitrarily, as shown in Fig. 1. We now partition the boundary into $m$ paths based on the common vertices,
$\partial A=P^{v_{1},v_{2}}_{\partial A}\bigcup P^{v_{2},v_{3}}_{\partial A}\bigcup\cdots\bigcup P^{v_{m-1},v_{m}}_{\partial A}\bigcup P^{v_{m},v_{1}}_{\partial A},$ (1)
where $P^{v_{i},v_{i+1}}_{{\partial A}}$ is a path on the boundary that includes edges between the vertex $v_{i}$ and the vertex $v_{i+1}$. Because $T_{A}$ is the spanning tree, there is a unique path between the vertex $v_{i}$ and the vertex $v_{i+1}$, which is denoted as $P^{v_{i},v_{i+1}}_{A}$. The union of $P^{v_{i},v_{i+1}}_{A}$ for all values of $i$ forms the spanning tree $T_{A}$,
$T_{A}=P^{v_{1},v_{2}}_{A}\bigcup P^{v_{2},v_{3}}_{A}\bigcup\cdots\bigcup P^{v_{m-1},v_{m}}_{A}\bigcup P^{v_{m},v_{1}}_{A}.$ (2)
Since one of the edges on the boundary $\partial A$ has been deleted, there are $m-1$ pairs of $(P^{v_{i},v_{i+1}}_{{\partial A}},P^{v_{i},v_{i+1}}_{A})$ that form joint cycles. Define the sum of the qubit values in the path $P^{v_{i},v_{i+1}}_{{\partial A}}$ as $\Sigma^{v_{i},v_{i+1}}_{\partial A}$, and that in the path $P^{v_{i},v_{i+1}}_{A}$ as $\Sigma^{v_{i},v_{i+1}}_{A}$. These joint cycles give rise to $m-1$ constraints, which can be expressed as
$\Sigma^{v_{i},v_{i+1}}_{\partial A}=\Sigma^{v_{i},v_{i+1}}_{A}=C_{\alpha}^{\,i}$ (3)
for all possible values of $i$, and $C_{\alpha}^{\,i}=0$ or $1$.
@@
For a given set of $\{C_{\alpha}^{\,i}\}$ (with $\alpha$ fixed), the $m-1$ constraints given by Eq. (3) impose $m-1$ constraints on the qubit values in the spanning tree of subsystem $A$. Since there are $|E_{T_{A}}|$ qubits in the spanning tree of subsystem $A$, the remaining $|E_{T_{A}}|-m+1$ qubits are independent. The qubit configurations of subsystem $A$ span a subspace $\Lambda_{\alpha}^{A}$ with dimension $d^{A}_{\alpha}=2^{|E_{T_{A}}|-m+1}$. The state of subsystem $A$ is then an equal superposition of all these allowable qubit configurations,
|$\ket{\psi_{\alpha}}=\frac{1}{\sqrt{d^{A}_{\alpha}}}\sum_{z\in\Lambda_{\alpha}^{A}}\ket{z}=\frac{1}{\sqrt{d^{A}_{\alpha}}}\sum_{z\in\Lambda_{\alpha}^{A}}\ket{z_{1},z_{2},\cdots,z_{n_{A}}}.$ (4)
@@
Similarly, for a given set of $\{C_{\alpha}^{\,i}\}$ (with $\alpha$ fixed), the $m-1$ constraints given by Eq. (3) impose $m-1$ constraints on the qubit values on the boundary $\partial A$. Since there are $|E_{\partial A}|$ qubits on the boundary, the remaining $|E_{\partial A}|-m$ qubits are independent. The qubit configurations of the boundary $\partial A$ span a subspace $\Lambda_{\alpha}^{B}$ with dimension $d^{B}_{\alpha}=2^{|E_{\partial A}|-m}$. The pair of subspaces $(\Lambda^{A}_{\alpha},\Lambda^{B}_{\alpha})$ is uniquely determined by the set $\{C_{\alpha}^{\,i}\}$ with $\alpha$ fixed.
@@
For a different set of $\{C_{\alpha^{\prime}}^{\,i}\}$, it follows from Eq. (3) that the subspaces $\Lambda^{B}_{\alpha}$ and $\Lambda^{B}_{\alpha^{\prime}}$ do not overlap and are therefore orthogonal. Since there are $|E_{\partial A}|-1$ independent qubits in the boundary $\partial A$ before joining the spanning trees $T_{A}$ and $T_{B}$, and the subspaces $\Lambda^{B}_{\alpha}$ and $\Lambda^{B}_{\alpha^{\prime}}$ have the same size, the number of distinct pairs of subspaces $(\Lambda^{A}_{\alpha},\Lambda^{B}_{\alpha})$ is $n_{\Lambda}=2^{m-1}$. After tracing out the degrees of freedom of subsystem $B$, we obtain the density matrix of subsystem $A$,
$\rho_{A}=\frac{1}{n_{\Lambda}}\sum_{\alpha=1}^{n_{\Lambda}}\ket{\psi_{\alpha}}\bra{\psi_{\alpha}}.$ | . (5)
@@
$S_{A}=-\log_{2}(n_{\Lambda})=m-1.$ (6)
The term $m$ is associated with the cycles formed by joining the spanning tree $T_{A}$ and the boundary $\partial A$, and depends on the size of the boundary $\partial A$, indicating its local nature; while the term $-1$ is associated with the cycle surrounding subsystem $A$ and is independent of the subsystem size, indicating a pattern of long-range entanglement.
@@
The entanglement entropy for the case where the boundary encompasses multiple cycles can also be evaluated (see Appendix B). We explicitly calculate the von Neumann entropy for several types of subsystem within a toric code (see Appendix C for details).
@@
General formalism for CSS codes. A general CSS code Calderbank and Shor (1996); Steane (1996) cannot always be represented by a simple graph like toric and surface codes. However, we can still leverage related concepts in graph theory (such as cycles and spanning trees) and generalize the above graph-based method to calculate the density matrix and von Neumann entropy of a subsystem in a CSS code, using its parity check matrix $H_{Z}$. This is because the multiplication of stabilizer generators can be represented by the addition (mod 2) of row vectors in the parity check matrix $H_{Z}$, and the latter is equivalent to the symmetric difference in the cycle space (see Appendix A for details). The $Z$-type stabilizer generators of CSS codes can be partitioned into three sets: (1) $s_{A}$ generators acting only on qubits within subsystem $A$, denoted as $\big{\{}S_{A}^{(i)}\big{\}}_{i=1}^{s_{A}}$;
(2) $s_{B}$ generators acting only on qubits within subsystem $B$, denoted as $\big{\{}S_{B}^{(i)}\big{\}}_{i=1}^{s_{B}}$;
and (3) $s_{AB}$ generators acting on qubits in both subsystems $A$ and $B$, denoted as $\big{\{}S_{AB}^{(i)}\big{\}}_{i=1}^{s_{AB}}$.
@@
For a specific stabilizer generator $S_{A}^{(i)}$ that acts only on qubits in subsystem $A$, one can delete one of its qubits to remove the cycle. This can be accomplished by first transforming the parity check matrix $H_{Z}$ into a specific form and then deleting the qubit and the stabilizer generator. Suppose we want to delete one of the qubits that the stabilizer $S_{A}^{(i)}$ acts on. The element of the parity check matrix corresponding to the row of $S_{A}^{(i)}$ and the column of this qubit is one. If another stabilizer generator, say $S$, also acts on this qubit, meaning that its element in the corresponding column is also one, we then replace the stabilizer generator $S$ with a new stabilizer generator $S^{\prime}=S_{A}^{(i)}S$, for which the element in the corresponding column becomes zero. We continue this process to replace all other stabilizer generators acting on this qubit with new stabilizer generators. The parity check matrix is then transformed into a form in which this qubit is acted on only by the stabilizer generator $S_{A}^{(i)}$. We delete the qubit and the stabilizer $S_{A}^{(i)}$ from the parity check matrix. The remaining matrix is still a valid parity check matrix, with its rows representing a different set of stabilizer generators. The procedure described above is equivalent to Gaussian elimination. We continue the above procedure to delete all stabilizer generators acting only on qubits in subsystem $A$ and those acting only on qubits in subsystem $B$. The parity matrix is transformed into a form like $M_{AB}=\begin{pmatrix}M_{A}&M_{B}\\ \end{pmatrix}$, which describes the stabilizer generators shared by subsystems $A$ and $B$.
@@
The boundary of subsystem $A$ may contain one or more cycles, which can be generated by multiplying several stabilizer generators shared between $A$ and $B$. This implies that new stabilizer generators can be defined from $M_{AB}$, which act only on qubits on the boundary of subsystem $A$. Similarly, if additional cycles exist within subsystem $A$, new stabilizer generators acting exclusively on qubits in $A$ can be defined from $M_{AB}$. Once all such generators have been identified, the Gaussian elimination procedure, along with appropriate row and column permutations, are applied to $M_{AB}$, transforming it into $W_{AB}$ (see Appendix D for details),
The submatrices $\begin{pmatrix}\mathbb{I}&W_{A}^{\prime}&\boldsymbol{0}&\boldsymbol{0}\end{pmatrix}$ and $\begin{pmatrix}\boldsymbol{0}&\boldsymbol{0}&W_{B}^{\prime}&\mathbb{I}\end{pmatrix}$ represent stabilizer generators enclosed entirely in subsystem $A$ and its boundary $\partial A$, respectively. The submatrix $\begin{pmatrix}W_{A}&W_{B}\\ \end{pmatrix}$ represents stabilizer generators formed by joining the spanning trees of subsystems $A$ and $B$. By reintroducing the deleted qubits and stabilizer generators, the original parity check matrix $H_{Z}$ can be transformed into $\tilde{H}_{Z}$ through Gaussian elimination,
@@
Define the qubit configuration of subsystem $A$ as $|z_{A}\rangle=|z_{A}^{d},z_{A}^{t}\rangle$ and that of subsystem $B$ as $|z_{B}\rangle=|z_{B}^{t},z_{B}^{d}\rangle$,
Here, $|z_{A}^{d}\rangle$ and |$\ket{z_{A}^{t}}$ represent the qubit configurations of the deleted and remaining qubits in subsystem $A$, respectively; while $|z_{B}^{d}\rangle$ and |$\ket{z_{B}^{t}}$ denote the qubit configurations of the deleted and remaining qubits in subsystem $B$, respectively. The qubit configuration |$\ket{z}=\ket{z_{A},z_{B}}$ satisfies $\tilde{H}_{Z}\,z=0$, which implies
@@
Suppose that the number of remaining qubits in subsystem $A$ is $t_{A}$ and that in subsystem $B$ is $t_{B}$, and the rank of the submatrices $W_{A}$ and $W_{B}$ is $r$. For a given $C_{\alpha}$, there are $2^{t_{A}-r}$ qubit configurations $|z_{A}\rangle$ that satisfy the constraints in Eq. (A graph-based approach to entanglement entropy of quantum error correcting codes). Consequently, these allowable configurations form a subspace $\Lambda_{\alpha}^{A}$ with dimension $d_{\alpha}^{A}=2^{t_{A}-r}$. Similarly, there are $2^{t_{B}-r}$ qubit configurations |$\ket{z_{B}}$ for the same $C_{\alpha}$, which form a subspace ${\Lambda}_{\alpha}^{B}$ with dimension ${d}_{\alpha}^{B}=2^{t_{B}-r}$. Different binary vectors $C_{\alpha}$ define mutually orthogonal subspaces for subsystems $A$ and $B$, which implies that the number of pairs of orthogonal subspaces $(\Lambda_{\alpha}^{A},\Lambda_{\alpha}^{B})$ is $n_{\Lambda}=2^{r}$. After tracing out the degrees of freedom of subsystem $B$, the density matrix of subsystem $A$ is given in the same form as Eq. (5), with $n_{\Lambda}=2^{r}$.
One can now straightforwardly calculate the von Neumann entropy,

### v3

@@ Abstract
We develop a graph-based method to study the entanglement entropy of Calderbank-Shor-Steane quantum codes. This method offers a straightforward interpretation for the entanglement entropy of quantum error correcting codes through graph-theoretical concepts, shedding light on the origins of both the local and long-range entanglement. Furthermore, it inspires an efficient computational scheme for evaluating the entanglement entropy. We illustrate the method by calculating the von Neumann entropy of subsystems in toric codes and two types of quantum low-density-parity check codes, and by comparing the scaling behavior of the entanglement entropy with respect to the subsystem size. Our method provides a new perspective for understanding the entanglement structure in quantum many-body systems.
@@
Introduction. Entanglement plays a crucial role in understanding the physics of quantum many-body systems, including topological matters Kitaev and Preskill (2006); Levin and Wen (2006); Eisert et al. (2010), the black hole information paradox Hawking (1976); Page (1993), information scrambling in a thermalization process Hayden and Preskill (2007), the connection between geometry and entanglement in quantum gravity Almheiri et al. (2015); Pastawski et al. (2015), and symmetry breaking/restoration Ares et al. (2023, 2024); Liu et al. (2024). In particular, the scaling law of the entanglement entropy with respect to subsystem size reveals key properties of quantum systems, e.g., discriminating critical and non-critical phases Vidal et al. (2003); Latorre et al. (2003). Entanglement is introduced in quantum error correcting codes to encode logical information in a non-local manner and protect it from local errors Shor (1995). This is manifested in their notable feature of local indistinguishability Knill and Laflamme (1997); Bennett et al. (1996). There exists a trade-off between the amount of entanglement and the code distance Bravyi et al. (2025); Li et al. (2025): the lower bound of the geometric entanglement is proportional to the code distance, which implies that more entanglement is required to correct more errors.
@@
Given the unique role of entanglement in quantum error correcting codes, it is important to develop methods to calculate the amount of entanglement within quantum codes. A group-theoretical method to evaluate the entanglement entropy for stabilizer codes was proposed by Refs. Fattal et al. (2004); Hamma et al. (2005a, b). It was used to study the topological entanglement entropy in toric codes Flammia et al. (2009) and other topological codes Kargarian (2008), and the entanglement dynamics in many-body systems Nahum et al. (2017); Li et al. (2019). Recently, a graph-based method was developed to study higher-rank topological phases Ebisu (2023).
@@
In this paper, we develop a different graph-based method to calculate the explicit form of the reduced density matrix and the entanglement entropy of Calderbank-Shor-Steane (CSS) quantum codes. It provides a clear picture on the origin of the local entanglement, which indicates the area law, and the long-range topological entanglement in the toric codes. Furthermore, it inspires an efficient algorithm to evaluate the entanglement entropy for all CSS codes, e.g., quantum low-density-parity-check (qLDPC) codes Tillich and Zémor (2013); Kovalev and Pryadko (2013); Breuckmann and Terhal (2016); Panteleev and Kalachev (2021); Breuckmann and Eberhardt (2021a); Panteleev and Kalachev (2022); Leverrier and Zémor (2022); Higgott and Breuckmann (2023); Wang et al. (2023); Lin and Pryadko (2024); Breuckmann and Eberhardt (2021b). We demonstrate the graph-based method by calculating the entanglement entropy for the bivariate-bicycle codes Bravyi et al. (2024) and quasi-cyclic codes Hagiwara and Imai (2007), and studying their scaling law with respect to the size of the subsystem.
@@
Two-dimensional toric code. We briefly introduce some basic concepts in graph theory Diestel (2024) that are essential for the subsequent discussions. A graph $G(V,E)$ consists of a set of vertices $V$ and a set of edges $E$. An edge connecting two vertices $v_{i},v_{j}\in V$ is represented by the pair $(v_{i},v_{j})$. A path is a sequence of distinct vertices connected by edges, and a cycle is a path that starts and ends at the same vertex. All cycles in a graph $G$ form a cycle space, the basis of which is the set of independent cycles. The cyclomatic number of $G$ is the count of these independent cycles SI (2025). A connected graph is one in which there exists at least one path between any pair of vertices. A tree graph is a connected graph that contains no cycles. A spanning tree of a graph $G$ is a connected subgraph that includes all vertices without cycles.
@@
Two-dimensional toric codes can be represented by simple two-dimensional graphs, where qubits are placed on edges Kitaev (2003). The faces and vertices of the graph represent stabilizers of type $Z$ and $X$, respectively (see Fig. 1 (a)). There are two distinct sets of cycles in the graph, the contractible cycles and non-contractible cycles. The former corresponds to $Z$-type stabilizers, while the latter represents logical Pauli $Z$ operators. Since all stabilizers have an eigenvalue of one in the code subspace, the sum of qubit values along any contractible cycle is zero (mod 2), imposing a constraint on these qubit values. The sum of qubit values along a non-contractible cycle can be either 0 or 1, depending on the specific logical state. If one of the logical Pauli $Z$ operators takes a specific eigenvalue, the qubit values along the corresponding non-contractible cycle are also subject to a constraint.
@@
We are concerned with the reduced density matrix and the von Neumann entropy of a subsystem $A$, a subset of qubits, of a toric code. The remaining qubits of the toric code are in the complementary subsystem $B$, as shown in Fig. 1 (b). There are cycles that are entirely enclosed by subsystem $A$ or subsystem $B$, as well as cycles shared by both subsystems, which are referred to as boundary stabilizers. The exterior boundary of subsystem $A$, denoted as $\partial A$, is defined as the set of edges (qubits) that belong to the boundary stabilizers but are not part of subsystem $A$.
@@
Figure 1: Illustration of subsystems and their spanning trees in a toric code embedded in a square lattice with periodic boundary conditions, initialized in the logical $Z$ eigenstate. (a) A $Z$-type stabilizer generator $S_{Z}$ acts on four qubits surrounding a face, while an $X$-type stabilizer generator $S_{X}$ acts on four qubits incident to a vertex. (b) Subsystem AA consists of qubits on blue edges and contains one cycle, highlighted by the orange face. Subsystem $B$ includes all remaining qubits on black edges. The exterior boundary of subsystem $A$ is indicated by the deep black edges. Stabilizers shared by subsystems $A$ and BB are marked by gray faces. (c) The spanning tree of subsystem $A$ is formed by removing one qubit (blue cross) from the orange face. The spanning tree of subsystem $B$ is constructed by removing qubits indicated by black crosses. (d) Blue edges $E_{T_{A}}$ indicate the spanning tree of subsystem $A$ and purple edges $E_{T_{B}}$ mark the spanning tree of subsystem $B$. The gray faces represent the joint cycles and the red vertices labeled by $v_{i}$ denote the common vertices of the two spanning trees. The entanglement entropy equals 55, which is the number of independent joint cycles, as determined by $|\mathcal{C}_{T_{A}\cup T_{B}}|$ or computed via Eq. (5).
Denote by $\mathcal{C}(A)$ and $\mathcal{C}(B)$ the sets of independent cycles in subsystems $A$ and $B$. By removing one edge from every independent cycle, we obtain spanning trees $T_{A}$ and $T_{B}$ within subsystems $A$ and $B$, respectively (see Fig. 1 (c)). The union of these two spanning trees $T_{A}$ and $T_{B}$ forms a graph $T_{A}\cup T_{B}$, which contains a set of independent joint cycles $\mathcal{C}(T_{A}\cup T_{B})$ (see Fig. 1 (d)). The sum of qubit values along any independent joint cycle $c_{k}\in\mathcal{C}(T_{A}\cup T_{B})$ is zero (mode 2), which implies
$\sum_{e\in E_{c_{k}}^{A}}z_{e}=\sum_{e\in E_{c_{k}}^{B}}z_{e}=C_{\alpha}^{k}~~(\text{mod}\,\,2).$ (1)
Here, $z_{e}=(1-\lambda)/2\in\{0,1\}$ is the qubit value on the edge $e$, with $\lambda=\pm 1$ the eigenvalue of Pauli $Z$ operator; and $E_{c_{k}}^{A}$ and $E_{c_{k}}^{B}$ represent the edge (qubit) sets of the partitions in the two subsystems; and $C_{\alpha}^{k}\in\left\{0,1\right\}$, with $\alpha$ labeling a bit string that represents the values of the sums in Eq. (1).
@@
For a given set of $\left\{C_{\alpha}^{k}\right\}$ (with $\alpha$ fixed), the $|\mathcal{C}(T_{A}\cup T_{B})|$ constraints given by Eq. (1) impose $|\mathcal{C}(T_{A}\cup T_{B})|$ constraints on the qubit values in the spanning tree of subsystem $A$. Since there are $|E_{T_{A}}|$ qubits in the spanning tree $T_{A}$, the remaining $|E_{T_{A}}|-|\mathcal{C}(T_{A}\cup T_{B})|$ qubits are independent. The qubit configurations of subsystem $A$ span a subspace $\Lambda^{A}_{\alpha}$ with dimension $d^{A}_{\alpha}=2^{|E_{T_{A}}|-|\mathcal{C}(T_{A}\cup T_{B})|}$. The state of the subsystem $A$ is then an equal superposition of all allowable qubit configurations,
$|\psi_{\alpha}\rangle=\frac{1}{\sqrt{d^{A}_{\alpha}}}\sum_{z~\in\Lambda^{A}_{\alpha}}|z_{1},z_{2},\cdots,z_{|E_{A}|}\rangle.$ (2)
@@
Similarly, for a given set of $\left\{C_{\alpha}^{k}\right\}$ (with $\alpha$ fixed), the $|\mathcal{C}(T_{A}\cup T_{B})|$ constraints given by Eq. (1) impose $|\mathcal{C}(T_{A}\cup T_{B})|$ constraints on the qubits values on the spanning tree of the subsystem $B$. Since there are $|E_{T_{B}}|$ qubits in the spanning tree $T_{B}$, the remaining $|E_{T_{B}}|-|\mathcal{C}(T_{A}\cup T_{B})|$ qubits are independent. The qubit configurations of the spanning tree $T_{B}$ also span a subspace $\Lambda^{B}_{\alpha}$ with the dimension $d^{B}_{\alpha}=2^{|E_{T_{B}}|-|\mathcal{C}(T_{A}\cup T_{B})|}$. The pair of subspaces $(\Lambda^{A}_{\alpha},\Lambda^{B}_{\alpha})$ is uniquely determined by the set $\{C^{k}_{\alpha}\}$ with $\alpha$ fixed.
@@
For a different set of $\left\{C_{\alpha^{\prime}}^{k}\right\}$, it follows from Eq. (1) that the subspaces $\Lambda^{B}_{\alpha}$ and $\Lambda^{B}_{\alpha^{\prime}}$ do not overlap and are therefore orthogonal. Since there are $|E_{T_{B}}|-|\mathcal{C}(T_{A}\cup T_{B})|$ independent qubits in the spanning tree $T_{B}$, and the subspaces $\Lambda^{B}_{\alpha}$ and $\Lambda^{B}_{\alpha^{\prime}}$ are of the same size, the number of distinct pairs of subspaces $(\Lambda^{A}_{\alpha},\Lambda^{B}_{\alpha})$ is $n_{\Lambda}=2^{|\mathcal{C}(T_{A}\cup T_{B})|}$. After tracing out the degrees of freedom of subsystem $B$, we obtain the density matrix of subsystem $A$,
@@
$S_{A}=-\log_{2}(n_{\Lambda})=|\mathcal{C}(T_{A}\cup T_{B})|.$ (4)
This result highlights a direct graph-theoretic interpretation of entanglement: it quantifies the independent joint cycles formed by combining the spanning tree structures of separated subsystems.
The entanglement entropy can be expressed in terms of the number of connected components within subsystems $A$ and $B$, as well as the number of common vertices shared between them SI (2025).
$S_{A}=|\mathcal{C}(T_{A}\cup T_{B})|=|V_{A\cap B}|-K_{1}-K_{2}+1,$ (5)
where $A\cap B$ is the intersection between $A$ and $B$, with $V_{A\cap B}$ the set of their common vertices, and $K_{1}$ and $K_{2}$ are the numbers of connected components within $A$ and $B$, respectively. This implies that the entanglement entropy depends on the global connectivity between subsystems, as well as the fragmentation within each subsystem.
@@
When subsystem $A$ is a single connected component and its boundary consists of only one cycle, its von Neumann entropy is $S_{A}=m-1$, with $m$ the number of boundary stabilizers. The term 1-1 is associated with the cycle surrounding subsystem $A$ and is independent of the size of subsystem, indicating a pattern of long-range entanglement. The entanglement entropy can also be evaluated for the case where the boundary encompasses multiple cycles. We explicitly calculate the von Neumann entropy for several types of subsystem within a toric code SI (2025).
@@
General formalism for CSS codes. A general CSS code Calderbank and Shor (1996); Steane (1996) cannot always be represented by a simple graph like toric and surface codes. However, we can still leverage related concepts in graph theory (such as cycles and spanning trees) and generalize the above graph-based method to calculate the density matrix and von Neumann entropy of a subsystem in a CSS code, using its parity check matrix $H_{Z}$. This is because the multiplication of stabilizer generators can be represented by the addition (mod 2) of row vectors in the parity check matrix $H_{Z}$, and the latter is equivalent to the symmetric difference in the cycle space SI (2025). The $Z$-type stabilizer generators of CSS codes can be partitioned into three sets: (1) $s_{A}$ generators acting only on qubits within subsystem $A$, denoted as $\big\{S_{A}^{(i)}\big\}_{i=1}^{s_{A}}$; (2) $s_{B}$ generators acting only on qubits within subsystem $B$, denoted as $\big\{S_{B}^{(i)}\big\}_{i=1}^{s_{B}}$; and (3) $s_{AB}$ generators acting on qubits in both subsystems $A$ and $B$, denoted as $\big\{S_{AB}^{(i)}\big\}_{i=1}^{s_{AB}}$.
@@
For a specific stabilizer generator $S_{A}^{(i)}$ that acts only on qubits in subsystem $A$, one can delete one of its qubits to remove the cycle. Suppose that we want to delete one of the qubits that the stabilizer $S_{A}^{(i)}$ acts on. If another stabilizer generator, say $S$, also acts on this qubit, meaning that its element in the corresponding column is also one, we then replace the stabilizer generator $S$ with a new stabilizer generator $S^{\prime}=S_{A}^{(i)}S$, for which the element in the corresponding column becomes zero. We continue this process to replace all other stabilizer generators acting on this qubit with new stabilizer generators. The parity check matrix is then transformed into a form in which this qubit is acted on only by the stabilizer generator $S_{A}^{(i)}$. We delete the qubit and the stabilizer $S_{A}^{(i)}$ from the parity check matrix. The remaining matrix is still a valid parity check matrix, with its rows representing a different set of stabilizer generators. The procedure described above is equivalent to Gaussian elimination. We continue the above procedure to delete all stabilizer generators acting only on qubits in subsystem $A$ and those acting only on qubits in subsystem $B$. The parity matrix is transformed into a form like $M_{AB}=\begin{pmatrix}M_{A}&M_{B}\\ \end{pmatrix}$, which describes the stabilizer generators shared by subsystems $A$ and $B$.
@@
The boundary of subsystem $A$ may contain one or more cycles, which can be generated by multiplying several stabilizer generators shared between $A$ and $B$. This implies that new stabilizer generators can be defined from $M_{AB}$, which act only on qubits on the boundary of subsystem $A$. Similarly, if additional cycles exist within subsystem $A$, new stabilizer generators acting exclusively on qubits in $A$ can be defined from $M_{AB}$. Once all such generators have been identified, we apply the above procedure to remove those cycles SI (2025).
By reintroducing the deleted qubits and stabilizer generators, the original parity check matrix $H_{Z}$ can be transformed into $\tilde{H}_{Z}$ through Gaussian elimination, along with appropriate row and column permutations,
@@
[…]
@@
[…]